OpenStudy (anonymous):

how do i simplify √(z^11/81z)

OpenStudy (anonymous):

first off \[\frac{z^{11}}{81z}=\frac{z^{10}}{81}\]

OpenStudy (anonymous):

square root of 81 is 9 square root of z^10 is z^5 because 10/2 = 5

OpenStudy (anonymous):

so answer is \[\frac{z^5}{9}\]

OpenStudy (anonymous):

sweet. thank you so much